OpenStudy (anonymous):

A progression has a first term of 12 and a fifth term of 18. Find the 13th term if the progression is geometric.

OpenStudy (callisto):

xr xr xr xr 12 ---> a2--->a3--->a4--->18 So 12(r^4) =18 \[r=\sqrt[4]{1.5}\] T(13) = 12r^(13-1) 12r^12...

OpenStudy (callisto):

sorry T(13) = 12r^(13-1) =12r^12 = 12[(1.5)^(1/4)]^12 = 12(1.5)^3 = 40.5...

OpenStudy (callisto):

Hmm.. Do you understand?

OpenStudy (anonymous):

Ok. THank you. Yes, I do. But i got 40.64

OpenStudy (callisto):

can you show your workings?

OpenStudy (anonymous):

18=12r^4 1.5=r^4 1.5^1/4=r r= 1.107 An= 12(1.107)^12 An= 40.6

OpenStudy (callisto):

it's because you round the number for r...